# Ladder-Residual: Parallelism-Aware Architecture for Accelerating Large Model Inference with Communication Overlapping

### v4

@@ Abstract
Large language model inference is both memory-intensive and time-consuming, often requiring distributed algorithms to efficiently scale. Various model parallelism strategies are used in distributed training and inference to partition computation across multiple devices, reducing memory load and computation time. However, using model parallelism necessitates communication of information between GPUs, which has been a major bottleneck and limits the gains obtained by scaling up the number of devices. We introduce Ladder Residual, a simple architectural modification applicable to all residual-based models that enables straightforward overlapping that effectively hides the latency of communication. Our insight is that in addition to system optimizations, one can also redesign the model architecture to decouple communication from computation. While Ladder Residual can allow communication-computation decoupling in conventional parallelism patterns, we focus on Tensor Parallelism in this paper, which is particularly bottlenecked by its heavy communication. For a Transformer model with 70B parameters, applying Ladder Residual to all its layers can achieve 29% end-to-end wall clock speed up at inference time with TP sharding over 8 devices. We refer the resulting Transformer model as the Ladder Transformer. We train a 1B and 3B Ladder Transformer from scratch and observe comparable performance to a standard dense transformer baseline. We also show that it is possible to convert parts of the Llama-3.1 8B model to our Ladder Residual architecture with minimal accuracy degradation by only retraining for 3B tokens. We release our code for training  [^1] Training code: https://github.com/IBM/dolomite-engine and inference  [^2] Inference code: https://github.com/mayank31398/ladder-residual-inference for easier replication of experiments.
@@ 1 Introduction
Past works have attempted to overlap the communication latency of TP by overlapping computation and communication. Chang et al. (2024) write fused kernels for AllGather followed by matmul and matmul followed by ReduceScatter. They break down matmuls into tiles and try to hide the latency of communicating a matmul tile with the computation of subsequent tiles. Jangda et al. (2022) propose CoCoNet, a domain-specific language to express distributed machine learning workloads. They propose to generate efficient GPU kernels for computation and communication using a custom compiler for the DSL. This approach has limited applicability with existing frameworks like PyTorch (Paszke et al., 2019) and JAX (Frostig et al., 2018) since the user needs to be well-acquainted with the DSL to generate efficient GPU kernels. Moreover, with the breakneck pace of accelerator and interconnect changes, these low-level systems optimizations require a rewrite for every new generation of hardware. However, there is a fundamental limit to how much communication latency can be reduced on such works which do not change the underlying model architecture. Instead of pure hardware optimizations (e.g., larger NVLink domain connecting 36 or 72 Blackwell GPUs) or pure low-level software optimization (e.g., rewriting all matmuls to overlap with communication), we explore model architectural changes that would enable a reduction in communication latency while maintaining accuracy. This makes our approach quite simple to apply in practice using a high level machine learning framework like PyTorch (Paszke et al., 2019) or JAX (Frostig et al., 2018) without writing any low-level device code.
@@
Table 1: Inference speedup from applying Ladder Residual on a Transformer model. The test setup is 1024 prompt length, 512 generated tokens, batch size 4, and TP world size of 8 for 1B, 3B, 8B, 34B, 70B, 176B model and TP world size of 16 for 405B model (using two nodes). The speedup value is calculated by comparing Ladder Transformer with Standard Transformer’s inference throughput in tokens per second. We measure the speedup both with and without P2P communication (modern server nodes usually support P2P communication). Note that disabling P2P communication significantly increases the AllReduce latency.
@@
  • Pretraining from scratch: We train a 1B and a 3B parameter Ladder Transformer model with 100B tokens on the FineWeb-edu dataset (Lozhkov et al., 2024) and compared it with the standard transformer of the same size trained on the same amount of tokens. We find that the Ladder Transformer matches the performance of the standard Transformer model.
@@ 2 Tensor Parallelism Background
A Transformer with N layers needs to perform the AllReduce 2N times and this can account for 38% of the inference latency for a 70B model using TP world size of 8, even with NVLink interconnect. If P2P communication is disabled, AllReduce latency can account for over 50% of the end-to-end latency. Modern nodes connect GPUs via NVLink but usually have no more than 8 GPUs per node, due to limited PCIe lanes or power and cooling constraints in datacenters. There is a steep falloff in communication bandwidth and sharp increase in latency when the communication happens across nodes either over InfiniBand or Ethernet thus making scaling TP practically infeasible outside a node.
Algorithm 1 Ladder Transformer Layer with Tensor Parallelism. Note that the AsyncAllReduce (AAR) returns a handle which is passed to the next layer.
@@ 4.1 Training From Scratch
We train a 1B and 3B Ladder Transformer from scratch and compare its performance with an equally sized standard Transformer model. All our models are trained on 100B tokens of FineWeb-edu dataset (HuggingFaceFW, 2024) using the StarCoder tokenizer (Li et al., 2023b). We also compare our model with the Parallel Attention/MLP architecture (Chowdhery et al., 2022; Wang & Komatsuzaki, 2021) which parallelizes the computation of the attention and the MLP module. This effectively reduces the communication cost by 50% for the tensor parallel AllReduce in both the forward and backward computation.
@@ 4.1.1 Experimental details
We use DDP (Distributed Data Parallel) (Li et al., 2020) to train the 1B and HSDP (Hybrid Sharded Data Parallel) (Zhao et al., 2023; Rajbhandari et al., 2020) to train the 3B models. For HSDP, we shard the model within 1 node (equipped with 8x H100 GPUs) and replicate the model outside the node. We use mixed precision training (Micikevicius et al., 2018) in BF16 (Kalamkar et al., 2019) with gradient accumulation and gradient AllReduce/ReduceScatter in FP32 for training stability. We train all our models with 2048 context length with a batch size of 4M tokens in a batch. The models are trained with cosine scheduler with a warmup of 8B tokens to a peak learning rate of $3\times 10^{-4}$. The learning rate is then decayed over 92B tokens to $3\times 10^{-5}$.
@@ 4.1.2 Results
The full results can be found at Table 3. We find that at the 1B model scale, Ladder Transformer achieves performance similar to Standard Transformer while beating the Parallel Transformer. At 3B parameter scale, we find that Standard Transformer is better than Ladder Transformer model with 3.2% lower perplexity and 1.2 points of absolute difference in accuracy. The Parallel Transformer has almost the same performance as Ladder Transformer at the 3B scale.
@@
Standard-Transformer-1B 34.22 70.33 41.10 71.49 87.30 55.41 59.98 18.54
Parallel-Transformer-1B 30.46 67.97 40.35 71.16 87.40 55.17 58.75 18.95
Ladder-Transformer-1B 31.31 67.76 41.18 71.49 86.60 55.17 58.92 18.42
Standard-Transformer-3B 38.99 74.12 46.48 74.59 92.00 58.48 64.11 14.48
Parallel-Transformer-3B 38.48 73.02 45.55 73.67 90.00 57.46 63.03 14.96
Ladder-Transformer-3B 36.77 72.43 45.66 73.72 89.90 58.96 62.91 14.90
@@ 4.2.1 Experimental details
We convert a state-of-the-art open-source model, Llama-3.1-8B-Instruct into a hybrid Ladder Residual structure, by applying Ladder Residual to the upper half of the model (layers 16-32 for LLaMA-3.1-8B-Instruct). We call this variant Hybrid-Ladder-8B-16L in Table 4. We also experiment with more aggressive adaptation where we applied Ladder Residual to the layers 12-32 and we call this experiment Hybrid-Ladder-8B-20L. We conduct supervised fine-tuning (SFT) for the resulting model on the 7M subset and the Gen subset of the Infinity-Instruct dataset [^3] https://huggingface.co/datasets/BAAI/Infinity-Instruct, which contains 3B tokens. We train for 2 epochs with AdamW optimizer with a batch size of 32. We use $5\times 10^{-6}$ learning rate with 200 steps of linear warmup, followed by cosine annealing to the end. We use Axolotl  [^4] https://github.com/axolotl-ai-cloud/axolotl for our SFT experiments.
@@
Hybrid-Ladder-8B-16L-retrained 65.93 59.13 42.20 78.86 39.66 80.29 59.10 59.02 21.95 56.24
@@
As shown in Table 4, after adaptation, there is a huge performance drop mainly on generative tasks as the computation flow is messed up. But after light retraining, the hybrid Ladder Llama is able to reach the same level of performance with the original Llama. By applying Ladder Residual on the last 16 layers, we can obtain 21% end-to-end wall clock speed up for inference with TP world size of 8 and batch size of 1. Our results demonstrate the potential of Ladder Residual being a drop-in adaptation technique for an existing model to make it faster without sacrificing performance. We additionally experiment with applying Ladder Residual to the last 20 layers of Llama and found that it leads to a slight drop in performance. There is a chance that with longer adaptation, or smarter adaptation techniques like distillation or iterative training, we can obtain a Ladder-Llama with more layers adapted. We leave the further exploration to future work.
@@ 5 Related Work
Prior works have explored various alternative designs for Transformer, for example parallel attention and mlp (Chowdhery et al., 2022; Wang & Komatsuzaki, 2021), linear attention (Katharopoulos et al., 2020), Grouped Query Attention (Ainslie et al., 2023), Cross-Layer Attention (Brandon et al., 2024) to improve the training and inference efficiency. Concurrently, Kraken (Prabhakar et al., 2024) treats each shard of weight as independent networks and only passes information between them occasionally to eliminate communication. However, the trained model under this design can only be used under a preset degree of parallelization. DistriFusion (Li et al., 2024) notices a similar phenomenon in the diffusion model, where the input is similar in adjacent diffusion steps, and overlaps the synchronization of distributed patches by reusing the feature maps from the previous diffusion step at the current step. Some of these variants are more widely adopted than others, due to the degree of impact they have on performance and efficiency. Past works have also considered adapting an existing checkpoint to these efficient variants.  Ainslie et al. (2023) extracted grouped-query attention from a multi-head attention model, and  Wang et al. (2024) converted a Llama model to a Mamba model by retraining on 50B tokens to close the performance gap. Comparatively, our adaptation is much lighter (3B tokens), showing that the representation shift introduced by Ladder Residual is easier to recover.  Wang et al. (2024) considered converting a Llama model to a Mamba model and used distillation to retrain the converted model. Such training paradigms that specifically tune the model to align with the original model could further improve the Ladder Residual based models.
@@ 6 Conclusion
We introduce Ladder Residual, architectural modifications that allow overlapping communication with computationfor model parallelism. We show that when running Tensor Parallelism, Ladder Residual can achieve great speed up across various model sizes, batch sizes, and number of GPUs. When applying Ladder Residual to Llama-3.1 8B Instruct, we only need lightweight retraining to reach the same level of performance as the original model while being 21% faster, showing its potential to be a plug-in for any pretrained Transformer. We also trained a 1B and 3B Ladder Transformer from scratch, and found that they are comparable to the standard Transformer of the same size while achieving over 55% speedup. Given that such a simple architectural change can obviate the need for expensive interconnects while maintaining model quality, we hope that our method will inspire even closer co-design between model architecture and inference systems.
7 Acknowledgement
We thank Tianle Cai, Devendra Chaplot, Igor Babuschkin, Albert Gu for helpful discussions and feedback that improved this work.
@@ A.1 PyTorch Profiler Trace
Here we provide the trace generated by the PyTorch Profiler [^5] https://pytorch.org/tutorials/recipes/recipes/profiler_recipe.html in Figure 5

### v5

@@ Abstract
Large language model inference is both memory-intensive and time-consuming, often requiring distributed algorithms to efficiently scale. Various model parallelism strategies are used to partition computation across multiple devices, reducing memory load and computation time. However, using model parallelism requires communication of information between GPUs, which limits the gains obtained by scaling up the number of devices. We introduce Ladder Residual, a simple architectural modification applicable to all residual-based models that enables straightforward overlapping to hide the latency of communication. Our insight is that in addition to system optimizations, the model architecture can also be redesigned to decouple communication from computation. While Ladder Residual can allow communication-computation decoupling in conventional parallelism patterns, we focus on Tensor Parallelism in this paper, which is particularly bottlenecked by its heavy communication. For a Transformer model with 70B parameters, applying Ladder Residual to all its layers can achieve 29% end-to-end wall clock speedup at inference time with sharding over 8 devices. We train a 1.2B and 3.5B Ladder Residual based Transformer models from scratch and observe comparable performance to a standard dense transformer baseline. We also show that it is possible to convert parts of the Llama-3.1 8B model to our Ladder Residual architecture with minimal accuracy degradation by only retraining for 3B tokens.
@@ 1 Introduction
Past works have attempted to overlap the communication latency of TP by overlapping computation and communication. Chang et al. (2024) write fused kernels for AllGather followed by matmul and matmul followed by ReduceScatter. They break down matmuls into tiles hide the latency of communicating a tile with the computation of subsequent tiles. Jangda et al. (2022) propose CoCoNet, a domain-specific language to express distributed machine learning workloads. They propose to generate efficient GPU kernels for computation and communication using a custom compiler for the DSL. This approach has limited applicability with existing frameworks like PyTorch (Paszke et al., 2019) and JAX (Frostig et al., 2018) since the user needs to be well-acquainted with the DSL to generate efficient GPU kernels. Moreover, with the breakneck pace of accelerator and interconnect changes, these low-level systems optimizations require a rewrite for every new generation of hardware. However, there is a fundamental limit to how much communication latency can be reduced with these approaches which do not change the underlying model architecture. Instead of pure hardware optimizations (e.g., larger NVLink domain connecting 36 or 72 NVIDIA Blackwell GPUs) or pure low-level software optimization (e.g., rewriting all matmuls to overlap with communication), we explore model architectural changes that would enable a reduction in communication latency while maintaining accuracy. This makes our approach quite simple to apply in practice using a high level machine learning framework like PyTorch (Paszke et al., 2019) or JAX (Frostig et al., 2018) without writing any low-level device code.
@@
Table 1: Inference speedup from applying Ladder Residual on a Transformer model. The test setup is 1024 prompt length, 512 generated tokens, batch size 4, and TP world size of 8 for 1B, 3B, 8B, 34B, 70B, 176B model and TP world size of 16 for 405B model (using two nodes). The models are designed according to Llama model families (and Bloom-176B). The speedup value is calculated by comparing Ladder Transformer with Standard Transformer’s inference throughput in tokens per second. We measure the speedup both with and without P2P communication (modern server nodes usually support P2P communication). Note that disabling P2P communication significantly increases the AllReduce latency.
@@
  • Pretraining from scratch: We train a 1.2B and a 3.5B parameter Ladder Transformer model with 100B tokens on the FineWeb-edu dataset (Lozhkov et al., 2024) and compared it with the standard transformer of the same size trained on the same amount of tokens. We find that the Ladder Transformer matches the performance of the standard Transformer model.
@@ 2 Tensor Parallelism Background
A Transformer with N layers needs to perform the AllReduce 2N times and this can account for 38% of the inference latency for a 70B model using TP world size of 8, even with NVLink interconnect. If P2P communication is disabled, AllReduce latency can account for over 50% of the end-to-end latency. Modern nodes connect GPUs via NVLink but usually have no more than 8 GPUs per node, due to limited PCIe lanes, power density and cooling constraints in datacenters. There is a steep falloff in communication bandwidth and sharp increase in latency when the communication happens across nodes either over InfiniBand or Ethernet thus making scaling TP practically infeasible outside a node.
Algorithm 1 Ladder Transformer Layer with Tensor Parallelism. Note that the AsyncAllReduce (ARR) returns a handle which is passed to the next layer.
@@ 4.1 Training From Scratch
We train a 1.2B and 3.5B Ladder Transformer from scratch and compare its performance with an equally sized standard Transformer model. All our models are trained on 100B tokens of FineWeb-edu dataset (HuggingFaceFW, 2024) using the StarCoder tokenizer (Li et al., 2023b). We also compare our model with the Parallel Attention/MLP architecture (Chowdhery et al., 2022; Wang & Komatsuzaki, 2021) which parallelizes the computation of the attention and the MLP module. This effectively reduces the communication cost by 50% for the tensor parallel AllReduce in both the forward and backward computation.
@@ 4.1.1 Experimental details
We use DDP (Distributed Data Parallel) (Li et al., 2020) to train the 1.2B and HSDP (Hybrid Sharded Data Parallel) (Zhao et al., 2023; Rajbhandari et al., 2020) to train the 3.5B models. For HSDP, we shard the model within 1 node (equipped with 8x H100 GPUs) and replicate the model outside the node. We use mixed precision training (Micikevicius et al., 2018) in BF16 (Kalamkar et al., 2019) with gradient accumulation and gradient AllReduce/ReduceScatter in FP32 for training stability. We train all our models with 2048 context length with a batch size of 4M tokens in a batch. The models are trained with cosine scheduler with a warmup of 8B tokens to a peak learning rate of $3\times 10^{-4}$. The learning rate is then decayed over 92B tokens to $3\times 10^{-5}$.
@@ 4.1.2 Results
The full results can be found at Table 3. We find that at the 1.2B model scale, Ladder Transformer achieves performance similar to Standard Transformer while beating the Parallel Transformer. At 3.5B parameter scale, we find that Standard Transformer is better than Ladder Transformer model with 3.2% lower perplexity and 1.2 points of absolute difference in accuracy. The Parallel Transformer has almost the same performance as Ladder Transformer at the 3.5B scale.
@@
Standard-Transformer-1.2B 34.22 70.33 41.10 71.49 87.30 55.41 59.98 18.54
Parallel-Transformer-1.2B 30.46 67.97 40.35 71.16 87.40 55.17 58.75 18.95
Ladder-Transformer-1.2B 31.31 67.76 41.18 71.49 86.60 55.17 58.92 18.42
Standard-Transformer-3.5B 38.99 74.12 46.48 74.59 92.00 58.48 64.11 14.48
Parallel-Transformer-3.5B 38.48 73.02 45.55 73.67 90.00 57.46 63.03 14.96
Ladder-Transformer-3.5B 36.77 72.43 45.66 73.72 89.90 58.96 62.91 14.90
@@ 4.2.1 Experimental details
We convert a state-of-the-art open-source model, Llama-3.1-8B-Instruct into a hybrid Ladder Residual structure, by applying Ladder Residual to the upper half of the model (layers 16-32 for LLaMA-3.1-8B-Instruct). We call this variant Hybrid-Ladder-8B-16L in Table 4. We also experiment with more aggressive adaptation where we applied Ladder Residual to the layers 12-32 and we call this experiment Hybrid-Ladder-8B-20L. We conduct supervised fine-tuning (SFT) for the resulting model on the 7M subset and the Gen subset of the Infinity-Instruct dataset [^1] https://huggingface.co/datasets/BAAI/Infinity-Instruct, which contains 3B tokens. We train for 2 epochs with AdamW optimizer with a batch size of 32. We use $5\times 10^{-6}$ learning rate with 200 steps of linear warmup, followed by cosine annealing to the end. We use Axolotl [^2] https://github.com/axolotl-ai-cloud/axolotl for our SFT experiments and Open LM Engine [^3] https://github.com/open-lm-engine/lm-engine for our pretraining experiments.
@@
Hybrid-Ladder-8B-16L-retrained 67.33 59.98 45.00 79.05 37.58 86.81 60.51 59.76 22.43 57.61
@@
As shown in Table 4, after adaptation, there is a huge performance drop mainly on generative tasks as the computation flow is messed up. But after light retraining, the hybrid Ladder Llama is able to reach the same level of performance with the original Llama. By applying Ladder Residual on the last 16 layers, we can obtain 21% end-to-end wall clock speed up for inference with TP world size of 8 and batch size of 1. Our results demonstrate the potential of Ladder Residual being a drop-in adaptation technique to make the model faster without sacrificing performance. We additionally experiment with applying Ladder Residual to the last 20 layers of Llama and found that it leads to a slight drop in performance. There is a chance that with longer adaptation, or smarter adaptation techniques like distillation or iterative training, we can obtain a Ladder-Llama with more layers adapted. We leave the further exploration to future work.
5 Discussion
5.1 Compatibility with other Parallelism Techniques
It should be noted that Ladder Residual is fully compatible with other model parallelism techniques like Pipeline Parallelism (PP). The AllReduce communication for TP is still asynchronous except at the PP boundary where the AllReduce needs to be waited upon to complete. Once the AllReduce result is available, we can forward 3 tensors (residual, current_mlp_output and current_attention_output) to the next pipeline stage. It should be noted that this is still pretty cheap since generally the P2P communication during inference is latency bound and can be implemented easily using the batch_isend_irecv API in PyTorch. Our approach also seemlessly works with Distributed Data Parallel (DDP) and Fully Sharded Data Parallel (FSDP).
5.2 Comparison to a 30% larger Ladder Transformer
Since the 70B Ladder Tranformer model achieves 30% higher tokens/sec compared to the 70B standard transformer as shown in Table 2, we compare the standard transformer to a 30% larger ladder transformer in Table 5. We find that the 30% larger ladder transformer is better than the standard transformer at both the 1.2B and 3.5B scale on average accuracy across bechmarks and wikitext perplexity while still achieving a much higher inference throughput.
Table 5: Performance of the Standard Transformer compared to a 30% larger Ladder Transformer at 1.2B and 3.5B scale, trained on FineWeb-edu for 100B tokens.
Model ARC-C ARC-E HellaSwag PIQA SciQ Winogrande Average Wikitext PPL Tokens/sec
Standard-Transformer-1.2B 34.22 70.33 41.10 71.49 87.30 55.41 59.98 18.54 1008.29
Ladder-Transformer-1.5B 33.96 70.16 42.58 71.98 87.90 55.41 60.33 17.47 1277.66
Standard-Transformer-3.5B 38.99 74.12 46.48 74.59 92.00 58.48 64.11 14.48 949.6
Ladder-Transformer-4.5B 40.96 75.00 46.81 73.99 90.80 57.70 64.21 14.05 1217.71
@@ 6 Related Work
Prior works have explored various alternative designs for Transformer, for example parallel attention and mlp (Chowdhery et al., 2022; Wang & Komatsuzaki, 2021), linear attention (Katharopoulos et al., 2020), Grouped Query Attention (Ainslie et al., 2023), Cross-Layer Attention (Brandon et al., 2024) to improve the training and inference efficiency. Some of these variants are more widely adopted than others, due to the degree of impact they have on performance and efficiency. Past works have also considered adapting an existing checkpoint to these efficient variants.  Ainslie et al. (2023) extracted grouped-query attention from a multi-head attention model, and  Wang et al. (2024) converted a Llama model to a Mamba model by retraining on 50B tokens to close the performance gap. Comparatively, our adaptation is much lighter (3B tokens), showing that the representation shift introduced by Ladder Residual is easier to recover.  Wang et al. (2024) considered converting a Llama model to a Mamba model and used distillation to retrain the converted model. Such training paradigms that specifically tune the model to align with the original model could further improve the Ladder Residual based models.
@@ 7 Conclusion
We introduce Ladder Residual, architectural modifications that allow overlapping communication with computationfor model parallelism. We show that when running Tensor Parallelism, Ladder Residual can achieve great speed up across various model sizes, batch sizes, and number of GPUs. When applying Ladder Residual to Llama-3.1 8B Instruct, we only need lightweight retraining to reach the same level of performance as the original model while being 21% faster, showing its potential to be a plug-in for any pretrained Transformer. We also trained a 1.2B and 3.5B Ladder Transformer from scratch, and found that they are comparable to the standard Transformer of the same size while achieving over 55% speedup. Given that such a simple architectural change can obviate the need for expensive interconnects while maintaining model quality, we hope that our method will inspire even closer co-design between model architecture and inference systems.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 PyTorch Profiler Trace
Here we provide the trace generated by the PyTorch Profiler [^4] https://pytorch.org/tutorials/recipes/recipes/profiler_recipe.html in Figure 5.